OpenStudy (anonymous):

{(x,y): x-3y=4} x-3y=4 Qx-6y=8 Find Q? It doesn't make sense to me.

OpenStudy (lyrae):

You have to give Q such a value it makes the equations in the system become equal x - 3y = 4 -> 2x - 6y = 8 | Qx - 6y = 8 Q = 2 We could similarly do Qx - 6y = 8 -> Q/2 x - 3y = 4 | 1 x - 3y = 4 Q/2 = 1 <=> Q = 2